# Characterizing NGC 6383: A study of pre-main sequence stars, mass segregation, and age using Gaia DR3 and 2MASS

### v1

@@ Abstract
This study employs Bayesian analysis and machine learning techniques to analyze the young open cluster NGC 6383 using data from Gaia DR3 and 2MASS. We identified 254 probable cluster members through HDBSCAN based on proper motions and determined the cluster’s core and tidal radius. To perform this analysis, we utilized an extension of Hamiltonian Monte Carlo, the No-U-Turn Sampler, using the Bayesian library PyMC. Our results indicate a mean cluster age of $4\pm 1$ Myr and a distance of $1.107\pm 0.035$ kpc. The analysis reveals mass segregation, especially among binary stars, and suggests that NGC 6383 is not fully relaxed. We observed a well-defined main sequence and a population of pre-main sequence stars, with a star formation range from $\sim 1$ to $6$ Myr, indicating recent star formation activity.
keywords:
open clusters and associations: individual — galaxies: star clusters: general — stars: distances — techniques: photometric — parallaxes — proper motions
pubyear: 2024pagerange: Characterizing NGC 6383: A study of pre-main sequence stars, mass segregation, and age using Gaia DR3 and 2MASSA
@@ 1 Introduction
Precise cluster’s members identification is a crucial step when it comes to a correct identification and characterization of the cluster. Various methods have been used for this process (Trumpler, 1930; Sanford, 1949; Aidelman et al., 2018; Fitzgerald et al., 1978; Pandey et al., 1989; Battinelli & Capuzzo-Dolcetta, 1991; Kharchenko et al., 2005; Paunzen et al., 2007; Lindoff, 1968), each offering different focuses and sometimes leading to consensus, or discrepancies. Recently, Bayesian analysis implementation on the membership identification, offers a different method to compare or compliment with the ones mentioned.
NGC 6383  [^1] NGC 6383, also known as NGC 6374 in the New General Catalog (Rauw & De Becker, 2008) and classified as Collinder 334 and Collinder 335 in the Collinder catalog, was initially misclassified in the original Collinder catalog (Collinder, 1931). is a young open cluster situated in the Carina-Sagittarius arm, within the Sh 2-012 star formation region (Sharpless, 1959). It forms part of the larger Sirius OB1 association, along with NGC 6530 and NGC 6531 (Rauw & De Becker, 2008). The galactic coordinates of the cluster are $\ell=355.68^{\circ}$ and $b=0.05^{\circ}$ (Conrad et al., 2017).
The distance to the cluster has been estimated by various authors, with an upper limit of $2130~{}\mathrm{pc}$ proposed by Trumpler (1930) and a lower limit of $760~{}\mathrm{pc}$ and $834~{}\mathrm{pc}$ suggested by Sanford (1949) and Aidelman et al. (2018), respectively.
@@
HD 159176, a double-lined spectroscopic binary composed of O7V type stars located in the projected center of the cluster, is responsible for ionizing the HII region (Penny et al., 2016). The estimated age of HD 159176 is around $2.3-2.8$ Myr (Rauw et al., 2010). A careful analysis of both the ages of the cluster and its central star system suggests that the central star predates the cluster itself. Furthermore, Fitzgerald et al. (1978) proposed that HD 159176 initiated star formation in the cluster’s core and beyond.
@@
The parameters of the cluster exhibit a wide range of results, contingent on the methodology utilized by the authors in their research. The cluster hosts sources that have recently embarked on the main sequence track, and there are also Pre-Main Sequence Stars (PMS) present Kalari (2019). Consequently, the age of the cluster has been a topic of considerable debate.
The implementation of Bayesian analysis and Machine Learning techniques in this study, explores a different method to obtained various parameters for characterizing NGC 6383 and contribute to the debate.
A summary of the historical results can be seen in Table 1.
@@
Trumpler (1930) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $2.13$cdots $\cdots$ $\cdots$
Sanford (1949) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $0.76$cdots $\cdots$ $\cdots$
Eggen (1961) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $0.30$cdots $\cdots$
Graham (1967) $74$cdots $\cdots$ $\cdots$ $1.38$cdots $10.68\pm 0.54$cdots
Evans (1968) At least $21$cdots $\cdots$ $\cdots$ $1.30$ $0.35$ $10.6$ $5$
Lindoff (1968) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $1.25$cdots $10.5$ $\sim 20$
Becker & Fenkart (1971) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $1.07$ $0.26$ $10.92$cdots
Fitzgerald et al. (1978) $\cdots$ $1.25$cdots $\cdots$ $1.50\pm 0.20$ $0.33\pm 0.02$ $10.85$ $1.7\pm 0.4$
Lloyd Evans (1978) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $1.35$ $0.35$ $10.6$ Up to $5$
The et al. (1985) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $1.40\pm 0.15$ $0.30\pm 0.01$cdots $\cdots$
Pandey et al. (1989) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $0.35$ $11.65$ $\sim 4.5$
Battinelli & Capuzzo-Dolcetta (1991) $27$cdots $\cdots$ $\cdots$ $1.38$ $0.34$cdots $\sim 20$
Feinstein (1994) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $1.40$ $0.33$cdots $\cdots$
Rauw et al. (2003) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $0.33$ $10.73$cdots
Kharchenko et al. (2005) $13$ $4.80$ $15$cdots $0.99$ $0.3$ $10.90$ $5$
Paunzen et al. (2007) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $1.7\pm 0.3$ $0.29\pm 0.05$cdots Less than $4$
Piskunov et al. (2008) $\cdots$ $\cdots$ $28.96\pm 6.63$cdots $0.99$ $0.30$ $10.90$ $5$
@@
Kalari (2019) At least $55$cdots $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $2.8\pm 1.6$
Jaehnig et al. (2021) $284$cdots $\cdots$ $0.93\pm 0.09$ $1.07$cdots $\cdots$ $\cdots$
He et al. (2022) $\cdots$ $\cdots$ $\cdots$ $0.88\pm 0.05$cdots $0.45$cdots $3.5$
Hunt & Reffert (2024) $322$ $2.49$ $22.73$ $0.89\pm 0.08$ $1.1$ $0.38$ $10.18$ $\sim 4$
Pulgar-Escobar & Henríquez-Salgado (2024) $266$ $1.21\pm 0.13$ $29.72\pm 7.69$ $0.92\pm 0.06$ $1.10\pm 0..04$ $0.47\pm 0.03$ $10.90\pm 0.33$ $\sim 1-4$
Table 1: Comprehensive historical review of observed parameters for the open cluster NGC 6383, gathered from various studies. The table details the number of cluster members, core and tidal radii (in arcminutes), parallax (in milliarcseconds), distance (in kiloparsecs), E(B-V) color excess, distance modulus $(m-M)$ (in magnitudes), and estimated ages (in millions of years). This summary provides a comparative analysis of how measurements and interpretations of NGC 6383 have evolved over time.
Figure 1: Membership probability of potential NGC 6383 members is plotted against their $G$ magnitudes, with colors representing astrometric fidelity. Each data point indicates a star, using a color gradient from blue to red to reflect the varying reliability of astrometric measurements. The horizontal dashed line at a probability of $0.6$ marks the threshold for probable cluster membership. As shown in the figure, stars with fainter magnitudes tend to have lower astrometric fidelity, which becomes more evident for stars with magnitudes fainter than 18.
Figure 2: Spatial distribution of potential members with a probability of at least 60 percent of NGC 6383 in Right Ascension $(\alpha)$ and Declination $(\delta)$, plotted against membership probability. Each point’s color indicates the probability of cluster membership, ranging from blue (lower probability) to red (higher probability). Concentric contour lines represent levels of Kernel Density Estimation (KDE) with an exponential kernel, revealing the density gradient of the points. The central coordinates of the cluster are marked by white dashed lines at $\alpha=263.6826^{\circ}$ and $\delta=-32.5838^{\circ}$, which correspond to the density peak within the plotted area. A scale bar of 5 parsecs is included at the bottom right to provide a reference for spatial scale.
2 Methodology
We acquired data from the Gaia third Data Release (DR3) (Gaia Collaboration et al., 2016, 2022), executing a cone search of $40~{}\mathrm{arcmin}$ radius on the Gaia archive, yielding $23740$ sources. We restricted the initial selection filtered sources to parallax ranges between $0.75$ and $1.1~{}\mathrm{mas}$—based on limits established in Pulgar-Escobar & Henríquez-Salgado (2024).
To enhance data reliability, we applied astrometric fidelity parameters from Rybizki et al. (2022). This parameter, derived from a neural network analysis of $17$ Gaia catalog metrics, assesses the astrometric solution’s trustworthiness. Sources with an astrometric fidelity above $0.5$ were retained, narrowing the selection to $15480$ sources, which is the $65$ per cent of the initial data set.
Subsequent refinement ensured inclusion of only sources with the position $(\alpha,\delta,\ell,b)$, proper motions $(\mu_{\alpha*},\mu_{\delta})$, parallax ($\varpi$) and the magnitudes $(G,G_{\mathrm{BP}},G_{\mathrm{RP}})$, finalizing the dataset at $15276$ sources. Systematic parallax offsets identified in Lindegren et al. (2021) were corrected using the Gaiadr3_zeropoint package. Moreover, to address the bias in proper motion for bright sources ($\mathrm{G}<13~{}\mathrm{mag}$), as discussed by Cantat-Gaudin & Brandt (2021), we applied a magnitude-based correction for sources with $\mathrm{G}=11-13~{}\mathrm{mag}$, compensating for up to $80~{}\mathrm{\mu as\,yr^{-1}}$ discrepancy between the frames of reference for bright and faint sources.
[…]
@@ 2.1 COSMIC
Characterization Of Star clusters using Machine learning Inference and Clustering (COSMIC) developed by Lucas Pulgar-Escobar et al. (in prep.), is a suite of functions designed for analyzing open clusters. Utilizing unsupervised machine learning algorithms, COSMIC processes extensive datasets, such as those from Gaia, to identify fundamental parameters of open clusters through clustering techniques and Bayesian estimation. As an open-source program, it is developed in Python 3.12 and integrates PyMC 5222www.pymc.io, a Python library specialized in Bayesian analysis.
The PyMC library aims to modernize and simplify Bayesian modeling. It offers a user-friendly interface for setting the prior distribution, generally a Normal distribution, and executes the sampling process efficiently. One of the key advantages of using PyMC is its implementation of the No-U-Turn Sampler (NUTS), an advanced MCMC algorithm that allows for efficient exploration of high-dimensional parameter spaces.
The NUTS algorithm builds upon Hamiltonian Monte Carlo (HMC) techniques, simulating a physical system to generate proposals. These proposals are based on the topology of the posterior distribution, allowing for an adaptive and balanced exploration of the parameter space. The algorithm terminates when a U-turn is detected, which serves as an indicator that the space has been sufficiently explored. A U-turn is detected when the trajectory of the sample path starts to double back on itself, indicating that it is moving in the opposite direction.
This efficiency and adaptability give NUTS an edge over traditional methods like Metropolis-Hastings and Gibbs sampling, which are often constrained by the shape of the distribution or the dimensions being considered (Phan et al., 2019).
Figure 3: Proper motions of sources with a probability of at least 60 percent of NGC 6383 in Right Ascension ($\mu_{\alpha}^{*}$) and Declination ($\mu_{\delta}$), color-coded by membership probability. The diagram uses hues transitioning from blue (lower probabilities) to red (higher probabilities), overlaid with density contours of proper motions. The contours illustrate the density gradient of proper motions. The center of proper motion distribution is marked by a blue cross, indicating the inferred mean proper motion values $(\mu_{\alpha}^{*},\mu_{\delta})=(2.5418,-1.7126)~{}\mathrm{mas}\,\mathrm{yr}^{-1}$, with corresponding orange dashed lines highlighting these central coordinates.

### v2

@@ Abstract
Context. This study focuses on the young open cluster NGC 6383, situated in the Carina-Sagittarius arm within the Sh 2-012 star formation region. Previous studies have provided estimates of the cluster’s distance, age, and structural properties. The presence of pre-main-sequence and low-mass stars, combined with the new data from Gaia DR3 and with 2MASS, significantly enhances the relevance of studying this cluster.
Aims. We aim to accurately identify cluster members, determine fundamental parameters, assess mass segregation, and establish precise age and distance using Gaia DR3 and 2MASS data.
Methods. We employed Bayesian analysis and machine learning techniques, including the Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) for member identification, the No-U-Turn Sampler (NUTS) from PyMC for modeling, the Sagitta neural network for the identification and age estimation of pre-main sequence stars, and ASteCA for isochrone fitting.
Results. We identified 254 probable cluster members with a mode cluster age of $3.53^{+1.40}_{-1.00}$ Myr and a distance of $1.11\pm 0.06$ kpc. The core and tidal radius were determined to be $1.95\pm 0.19$ and $40.4\pm 14.3\,\mathrm{arcmin}$, respectively. The analysis revealed primordial mass segregation among binary stars, indicating that NGC 6383 is not fully relaxed. The color magnitude diagram (CMD) shows a well-defined main sequence and a population of pre-main sequence stars, suggesting recent star formation activity from approximately 1 to 6 Myr ago. The software used for this investigation is released with the paper.
Conclusions. Our analysis provides updated parameters for NGC 6383, confirming relative recent star formation and mass segregation, and demonstrating the effectiveness of combining advanced computational techniques with traditional methods for studying stellar clusters.
Key Words.:
open clusters and associations: individual – galaxies: star clusters: general – stars: distances – techniques: photometric – parallaxes – proper motions
@@ 1 Introduction
The estimation of membership is a crucial step in the identification and characterization of star clusters. Various methods have been proposed in Trumpler (1930); Sanford (1949); Aidelman et al. (2018); Fitzgerald et al. (1978); Pandey et al. (1989); Battinelli & Capuzzo-Dolcetta (1991); Kharchenko et al. (2005); Paunzen et al. (2007); Lindoff (1968), each focusing on certain properties and sometimes leading to consensus or discrepancies. Recently, the implementation of Bayesian analysis for membership identification provides a different method to compare or complement the ones mentioned.
@@
HD 159176, a double-lined spectroscopic binary composed of O7V type stars located in the projected center of the cluster, is responsible for ionizing the HII region (Penny et al., 2016). The estimated age of HD 159176 is $2.30-2.80$ Myr (Rauw et al., 2010). A careful analysis of both the ages of the cluster and its central star system suggests that the central star predates the cluster itself. Furthermore, Fitzgerald et al. (1978) proposed that HD 159176 initiated star formation in the cluster’s core and beyond.
@@
This paper is structured as follows: Section 2 outlines the methodology, covering data acquisition and processing methods, including the use of the Gaia DR3 and 2MASS catalogs, and describes the membership determination using HDBSCAN and the Bayesian analysis for parameter estimation. Section 3 presents the results, including the determination of the cluster’s age, distance, and structural properties, the analysis of mass segregation, and the color-magnitude diagram. It also discusses the findings in the context of previous studies, addressing peculiar stars and comparisons with historical data. Finally, Section 4 concludes the study, summarizing the key results and their implications. The photometric system adopted includes the Gaia DR3 photometric bands ($G,G_{BP},G_{RP}$) and the 2MASS bands ($J,H,K_{s}$).
Figure 1: Membership probability of potential NGC 6383 members (membership probability $>50\%$) is plotted against their $G$ magnitudes, with colors representing astrometric fidelity. Each data point indicates a star, using a color gradient from blue to red to reflect the varying reliability of astrometric measurements. The horizontal dashed line at a probability of $0.6$ marks the threshold used to distinguish between probable members (with membership probabilities between $60\%$ and $80\%$) and members (with membership probabilities $\geq 80\%$) of the cluster. As shown in the figure, stars with fainter magnitudes tend to have lower astrometric fidelity. This becomes more evident at $G>18\leavevmode\nobreak\ \mathrm{mag}$.
2 Methodology
We acquired data from the Gaia third Data Release (DR3) (Gaia Collaboration et al., 2016, 2023), executing a cone search of $40.0\leavevmode\nobreak\ \mathrm{arcmin}$ radius in the Gaia archive, which yielded $23740$ sources. We restricted the initial selection of filtered sources to parallax ranges between $0.750$ and $1.10\leavevmode\nobreak\ \mathrm{mas}$-based on limits established in Pulgar-Escobar & Henríquez-Salgado (2024).
To enhance data reliability, we applied the astrometric fidelity parameter from Rybizki et al. (2022). This parameter, derived from a neural network analysis of 17 Gaia catalog metrics, assesses the trustworthiness of the astrometric solution. Sources with an astrometric fidelity above 0.5 were retained, narrowing the selection to 15480 sources, which represent 65 percent of the initial data set.
Figure 2: Spatial distribution of probable members and members of NGC 6383 in Right Ascension $(\alpha)$ and Declination $(\delta)$, color-coded according to their membership probability. Each color indicates the probability of cluster membership. The contours represent levels of Kernel Density Estimation (KDE) with an exponential kernel, revealing the density gradient of the points. The central coordinates of the cluster are marked by white dashed lines at $\alpha=263.683^{\circ}$ and $\delta=-32.584^{\circ}$, which correspond to the density peak within the plotted area. A scale bar of 5 parsecs is included at the bottom right to provide a reference for the spatial scale.
Subsequent refinement ensured the inclusion of only sources for which all parameters $(\alpha,\delta,\ell,b)$, proper motions $(\mu_{\alpha*},\mu_{\delta})$, parallax ($\varpi$), and magnitudes $(G,G_{\mathrm{BP}},G_{\mathrm{RP}})$ were available, reducing the dataset to $15276$ sources. Systematic parallax offsets identified in Lindegren et al. (2021) were corrected using the Gaiadr3_zeropoint package. Moreover, to address the bias in proper motion for bright sources ($\mathrm{G}<13\leavevmode\nobreak\ \mathrm{mag}$), as discussed in Cantat-Gaudin & Brandt (2021), we applied a magnitude-based correction for sources with $\mathrm{G}=11-13\,\mathrm{mag}$, compensating for up to $80\leavevmode\nobreak\ \mathrm{\mu as\,yr^{-1}}$ discrepancy between the frames of reference for bright and faint sources.
[…]
Table 1: Historical review of observed parameters for the open cluster NGC 6383, gathered from various studies. The table details the number of cluster members, core and tidal radii, parallax, distance, E(B-V) color excess, distance modulus $(m-M)$, and estimated ages. This summary provides a view on how measurements and interpretations of NGC 6383 have evolved over time.
@@ 2 Methodology
Trumpler (1930) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 2.13 $\cdots$ $\cdots$ $\cdots$
Sanford (1949) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 0.76 $\cdots$ $\cdots$ $\cdots$
Eggen (1961) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ 0.30 $\cdots$ $\cdots$
Graham (1967) 74 $\cdots$ $\cdots$ $\cdots$ 1.38 $\cdots$ 10.7$\pm$0.5 $\cdots$
Evans (1968) At least 21 $\cdots$ $\cdots$ $\cdots$ 1.30 0.35 10.6 5.0
Lindoff (1968) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 1.25 $\cdots$ 10.5 $\sim$ 20
Becker & Fenkart (1971) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 1.07 0.26 10.9 $\cdots$
Fitzgerald et al. (1978) $\cdots$ 1.25 $\cdots$ $\cdots$ 1.50$\pm$0.20 0.33$\pm$0.02 10.9 1.7$\pm$0.4
Lloyd Evans (1978) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 1.35 0.35 10.6 Up to 5.0
The et al. (1985) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 1.40$\pm$0.15 0.30$\pm$0.01 $\cdots$ $\cdots$
Pandey et al. (1989) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ 0.35 11.7 $\sim$ 4.5
Battinelli & Capuzzo-Dolcetta (1991) 27 $\cdots$ $\cdots$ $\cdots$ 1.38 0.34 $\cdots$ $\sim$ 20
Feinstein (1994) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 1.40 0.33 $\cdots$ $\cdots$
Rauw et al. (2003) $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ 0.33 10.7 $\cdots$
Kharchenko et al. (2005) 13 4.80 15.0 $\cdots$ 0.99 0.30 10.9 5.0
Paunzen et al. (2007) $\cdots$ $\cdots$ $\cdots$ $\cdots$ 1.70$\pm$0.30 0.29$\pm$0.05 $\cdots$ Less than 4.0
Piskunov et al. (2008) $\cdots$ $\cdots$ 29.0$\pm$6.6 $\cdots$ 0.99 0.30 10.9 5.0
@@
Kalari (2019) At least 55 $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ $\cdots$ 2.8$\pm$1.6
Jaehnig et al. (2021) 284 $\cdots$ $\cdots$ 0.93$\pm$0.09 1.07 $\cdots$ $\cdots$ $\cdots$
He et al. (2022) $\cdots$ $\cdots$ $\cdots$ 0.88$\pm$0.05 $\cdots$ 0.45 $\cdots$ 3.5
Hunt & Reffert (2024) 322 2.49 22.7 0.89$\pm$0.08 1.10 0.38 10.2 $\sim$ 4.0
Pulgar-Escobar & Henríquez-Salgado (2024) 266 1.21$\pm$0.13 29.7$\pm$7.7 0.92$\pm$0.06 1.10$\pm$0.04 0.47$\pm$0.03 10.9$\pm$0.33 $\sim$ 1-4
Figure 3: Proper motions of probable members and members of NGC 6383 in Right Ascension ($\mu_{\alpha}^{*}$) and Declination ($\mu_{\delta}$), the symbols are color-coded according to the membership probability of the stars. The diagram uses hues transitioning from blue (lower probabilities) to red (higher probabilities), overlaid with density contours of proper motions. The contours illustrate the density gradient of the proper motions. The center of proper motion distribution is marked with a blue cross, indicating the inferred mean proper motion values $(\mu_{\alpha}^{*},\mu_{\delta})=(2.54,-1.71)\leavevmode\nobreak\ \mathrm{mas}\,\mathrm{yr^{-1}}$, with corresponding orange dashed lines highlighting these central coordinates.
Figure 4: Left panel: Distributions of the parallax measurements in NGC 6383. U. parallax in red represents measurements with fractional parallax errors less than $0.1$, used for the parallax estimation, Parallax in orange represents all the data, and N.U. Parallax in green represents sources with a fractional parallax error greater than 0.100, not used for the parallax estimation. The blue curve shows a Gaussian fit to the U. parallax distribution, with the mean parallax value $\mu_{\varpi}=0.908\pm 0.004\,\mathrm{mas}$ marked with a dashed green line. Middle panel: G magnitude versus parallax values, with the sources utilized for the distance estimation in blue and discarded due to having a fractional parallax error greater than 0.100 in red. Gray error bars illustrate the uncertainty in parallax measurements, especially pronounced at fainter magnitudes. Right panel: Histograms of geometric distances derived from Bailer-Jones et al. (2021) are shown, distinguishing between used (red) and all measured (orange) parallaxes, including those not used (green). The central dashed blue line indicates the mode of the sampled distance ($1.1100\,\mathrm{kpc}$), with its mean standard deviation of $0.0599\,\mathrm{kpc}$ shaded in blue. For convenience, only the portion of the histogram corresponding to distances within $0.8$ to $1.2$ times the minimum and maximum used distances is displayed.
@@ 2.1 COSMIC
The Characterization Of Star clusters using Machine learning Inference and Clustering (COSMIC) developed by Lucas Pulgar-Escobar et al. (in prep.), is a suite of functions designed for analyzing open clusters. COSMIC utilizes unsupervised machine learning algorithms to process extensive datasets, such as those from Gaia, to identify fundamental parameters of open clusters through clustering techniques and Bayesian estimation. As an open-source program, it is developed in Python 3.12 and integrates PyMC 5222www.pymc.io, a Python library specialized in Bayesian analysis.
The PyMC library aims to modernize and simplify Bayesian modeling. It offers a user-friendly interface for setting prior distributions and executes the sampling process efficiently. One of the key advantages of using PyMC is its implementation of the No-U-Turn Sampler (NUTS), an advanced MCMC algorithm that allows for efficient exploration of high-dimensional parameter spaces.
The NUTS algorithm builds upon Hamiltonian Monte Carlo (HMC) techniques, simulating a physical system to generate proposals. These proposals are based on the topology of the posterior distribution, allowing for an adaptive and balanced exploration of the parameter space. The algorithm terminates when a U-turn is detected, which serves as an indicator that the space has been sufficiently explored. A U-turn in the trajectory of the MCMC algorithm is detected when the trajectory of the sample path starts to bend on itself, indicating that the sampler is proceeding in the opposite direction.
Its efficiency and adaptability give NUTS an edge over traditional methods like Metropolis-Hastings and Gibbs sampling, which are often constrained by the shape of the distribution or the dimensions being considered (Phan et al., 2019).